# Scale-invariant Phenomena in Repeating Fast Radio Bursts and Glitching Pulsars

### v1

@@ 3 Scale Invariance in Glitching Pulsars
Remarkably, both a glitch and an antiglitch accompanied by FRB-like bursts from SGR J1935+2154 were recently reported. The temporal coincidences between the glitch/antiglitch and FRB-like bursts imply some physical connection between them (Ge et al., 2022; Younes et al., 2023). The statistical distributions of glitch sizes ($\Delta\nu$) and times between successive glitches (waiting times) for some individual pulsars are power-law forms (Morley & Garcia-Pelayo, 1993; Melatos et al., 2008; Carlin & Melatos, 2021). Power-law distributions of energy and waiting time are also found in some repeating FRBs (Wang & Yu, 2017; Katz, 2018; Cheng et al., 2020; Zhang et al., 2021; Hewitt et al., 2022; Jahns et al., 2023; Wang et al., 2023). Inspired by the associations between the glitch/antiglitch and FRB-like bursts of SGR J1935+2154 and the statistical similarities between pulsar glitches and repeating FRBs, here we investigate whether pulsar glitches have a similar scale-invariant behavior.
Figure 4: Scale-invariant structure of glitches of PSR B1737–30. CDFs of fluctuations of glitch size (left panel) and waiting time (middle panel) for $n=1$ (red dots), $n=5$ (green dots), and $n=10$ (blue dots), and the best-fitting $q$-Gaussian distributions (solid curves). Right panel: the best-fitting $q$ values in the $q$-Gaussian distribution as a function of $n$.
@@
For the detected glitches in PSR B1737–30, the CDFs of fluctuations of glitch size and waiting time are presented in the left and middle panels of Figure 4, respectively. It is clear that the $q$-Gaussian function fits the data points very well. To better highlight the scale-invariant property in PSR B1737–30’ glitches, the best-fitting $q$ values as a function of $n$ for glitch size and waiting time are plotted in the right panel of Figure 4. Again, we find that the $q$ values are nearly constant and independent of $n$. We also list the average $q$ values from $n=1$ to 10 for glitch size ($\overline{q}=2.38\pm 0.04$) and waiting time ($\overline{q}=1.80\pm 0.18$) in Table 1. Interestingly, the average $q$ values of glitch size (energy) and waiting time are well consistent with those of FRB 20201124A’s high-energy bursts, which indicate that there may be an underlying physical connection between pulsar glitches and FRB 20201124A’s high-energy bursts.
@@ 4 Summary and discussion
SOC dynamics can be effectively identified and diagnosed by analyzing the avalanche size fluctuations (Caruso et al., 2007). When criticality appears, the PDFs for the avalanche size fluctuations at different times have fat tails with a $q$-Gaussian form. Such a $q$-Gaussian behavior is independent of the time interval adopted, and it is found so when considering energy fluctuations between real earthquakes (Caruso et al., 2007). That is, the $q$ values in $q$-Gaussian distributions are nearly invariant for different temporal interval scales, implying a scale-invariant structure in earthquakes (see also Wang et al. 2015). Analogous scale-invariant characteristics have also been discovered in some astronomical phenomena, such as SGRs (Chang et al., 2017; Wei et al., 2021; Sang & Lin, 2022), the repeating FRB 20121102 (Lin & Sang, 2020; Wei et al., 2021), and GRB X-ray flares (Wei, 2023) and precursors (Li & Yang, 2023). Given the temporal coincidences between the glitch/antiglitch and FRB-like bursts from the Galactic magnetar SGR J1935+2154 (Ge et al., 2022; Younes et al., 2023), here we examined their possible physical connection by comparing the scale-invariant similarities between repeating FRBs and glitching pulsars.
For repeating FRBs, we focused on the statistical properties of the hyperactive repeating source FRB 20201124A using two samples from different observations. The first sample contains 2610 bursts from the FAST observation, and the second one consists of 46 high-energy bursts from the observations of four small 25–32-m class radio bursts. Note that except for FRB 20121102, it is unclear whether other repeating FRBs share the same scale-invariant property. With the FAST data, we confirmed that the PDFs of fluctuations of energy, peak flux, duration, and waiting time can be well fitted by a $q$-Gaussian function, and the $q$ values keep nearly steady for different time scales. Moreover, we found that the average $q$ values of energy, peak flux, and duration for the FAST data of FRB 20201124A are very close to those of FRB 20121102 (Wei et al., 2021), which indicate that there is a common scale-invariant feature in repeating FRBs. With the high-energy burst sample, similar scale-invariant results were obtained, but the average $q$ values of duration and waiting time of high-energy bursts are significantly different from those of FAST observations of low-energy bursts. This implies that low- and high-energy bursts may originate from different emission mechanisms or emission regions at the progenitor source. Wang et al. (2015) also found that different faulting styles correspond to different $q$ values in earthquakes.
@@
This work is partially supported by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 12373053, 12321003, and 12041306), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the CAS Project for Young Scientists in Basic Research (grant No. YSBR-063), the CAS Organizational Scientific Research Platform for National Major Scientific and Technological Infrastructure: Cosmic Transients with FAST, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).

### v2

@@ 3 Scale Invariance in Glitching Pulsars
Figure 4: Statistical results of glitches of PSR B1737–30. Upper-left panel: The CDFs of glitch size fluctuations $X_{n}=S_{i+n}-S_{i}$ for different temporal interval scales $n=1$ (red dots), $n=5$ (green dots), and $n=10$ (blue dots). $X_{n}$ is rescaled by the standard deviation of $X_{n}$, i.e. $x_{n}=X_{n}/\sigma_{X_{n}}$. The data are fitted by the CDF of $q$-Gaussian function (solid curves). Upper-right panel: The same data process and the fitting to the CDF of $q$-Gaussian function, except now for the CDFs of waiting-time fluctuations. Lower panel
: The best-fitting $q$ values in the $q$-Gaussian distribution as a function of $n$.
Remarkably, both a glitch and an antiglitch accompanied by FRB-like bursts from SGR J1935+2154 were recently reported. The temporal coincidences between the glitch/antiglitch and FRB-like bursts imply some physical connection between them (Ge et al., 2022; Younes et al., 2023). The statistical distributions of glitch sizes ($\Delta\nu$) and times between successive glitches (waiting times) for some individual pulsars are power-law forms (Morley & Garcia-Pelayo, 1993; Melatos et al., 2008; Carlin & Melatos, 2021). Power-law distributions of energy and waiting time are also found in some repeating FRBs (Wang & Yu, 2017; Katz, 2018; Gourdji et al., 2019; Cheng et al., 2020; Cruces et al., 2021; Li et al., 2021b; Zhang et al., 2021; Hewitt et al., 2022; Jahns et al., 2023; Wang et al., 2023). Inspired by the associations between the glitch/antiglitch and FRB-like bursts of SGR J1935+2154 and the statistical similarities between pulsar glitches and repeating FRBs, here we investigate whether pulsar glitches have a similar scale-invariant behavior.
@@
[…]
@@ 4 Summary and discussion
SOC dynamics can be effectively identified and diagnosed by analyzing the avalanche size fluctuations (Caruso et al., 2007). When criticality appears, the PDFs for the avalanche size fluctuations at different times have fat tails with a $q$-Gaussian form. Such a $q$-Gaussian behavior is independent of the time interval adopted, and it is found so when considering energy fluctuations between real earthquakes (Caruso et al., 2007). That is, the $q$ values in $q$-Gaussian distributions are nearly invariant for different temporal interval scales, implying a scale-invariant structure in earthquakes (see also Wang et al. 2015). Analogous scale-invariant characteristics have also been discovered in some astronomical phenomena, such as SGRs (Chang et al., 2017; Wei et al., 2021; Sang & Lin, 2022), the repeating FRB 20121102 (Lin & Sang, 2020; Wei et al., 2021), GRB X-ray flares (Wei, 2023) and precursors (Li & Yang, 2023), and gamma-ray flares of the Sun and the blazar 3C 454.3 (Peng et al., 2023). Given the temporal coincidences between the glitch/antiglitch and FRB-like bursts from the Galactic magnetar SGR J1935+2154 (Ge et al., 2022; Younes et al., 2023), here we examined their possible physical connection by comparing the scale-invariant similarities between repeating FRBs and glitching pulsars.
For repeating FRBs, we focused on the statistical properties of the hyperactive repeating source FRB 20201124A using two samples from different observations. The first sample contains 2744 bursts from the FAST observation, and the second one consists of 46 high-energy bursts from the observations of four small 25–32-m class radio telescopes. Note that except for FRB 20121102, it is unclear whether other repeating FRBs share the same scale-invariant property. With the FAST data, we confirmed that the PDFs of fluctuations of energy, peak flux, duration, and waiting time can be well fitted by a $q$-Gaussian function, and the $q$ values keep nearly steady for different time scales. Moreover, we found that the average $q$ values of energy, peak flux, and duration for the FAST data of FRB 20201124A are very close to those of FRB 20121102 (Wei et al., 2021), which indicate that there is a common scale-invariant feature in repeating FRBs. With the high-energy burst sample, similar scale-invariant results were obtained, but the average $q$ values of duration and waiting time of high-energy bursts are significantly different from those of FAST observations of low-energy bursts. This implies that low- and high-energy bursts may originate from different emission mechanisms or emission regions at the progenitor source. Wang et al. (2015) also found that different faulting styles correspond to different $q$ values in earthquakes.
@@
We are grateful to the anonymous referee for constructive comments that helped improve our work. This work is partially supported by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 12373053, 12321003, and 12041306), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the CAS Project for Young Scientists in Basic Research (grant No. YSBR-063), the CAS Organizational Scientific Research Platform for National Major Scientific and Technological Infrastructure: Cosmic Transients with FAST, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).